two bodies of diferrent masses Ma and Mb are dropped from same  height  h .what is the ratio of time taken by the two bodies to fall through this ditance

Grade:11

1 Answers

Sanju
6 years ago
We have, s = ut+ 0.5 at
where, s = height, u = initial velocity, t= time taken and a = acclrn due to gravity
Here u = 0
So the eqn becomes s = 0.5 at2
Or, t = sqrt(2s/g)
This relation shows that time depends on the height only and is independent of the mass.
Therefore, ratio of time taken is:
t/ t= sqrt (s1 / s2)